Grade 11General Physics

wo stones are dropped from top of tower a half a second apart.the time after dropping the first stone at which the distance between two stones is 20m is

shyam sundar
14 Years agoGrade 11

2 Answers

abhishek upadhyay
ApprovedApproved Tutor Answer14 Years ago

if first block travels for t time then second travels for (t- 1/2)s

  h1=5t2                      taking g =10

h2 = 5 (t - 1/2)2       substracting

20 = 5(t - 1/4)

or t= 4.25s

xyz xz
ApprovedApproved Tutor Answer14 Years ago

As intial velocity is 0.

distance travelled by 1st stone for t seconds is S= 1/2 t2

distance travelled by 2nd stone is S= 1/2 (t-(1/2))2   { since 2nd body is throwned after 1/2 second}

and separation is 20 m

therefore;

                      [1/2 t2] - [ 1/2  (t-(1/2))2 ] = 20

                       implies t=4.25 seconds